OpenStudy (anonymous):

medal Complete the general form of the equation of a sinusoidal function having an amplitude of 6, a period of 2pie/3 , and a phase shift to the left 1 unit. y =

OpenStudy (chrisplusian):

\[6\sin(\frac{ 2\pi }{ 3 }+1)\]

OpenStudy (chrisplusian):

\[6\sin(\frac{ 2\pi }{ 3 }x+1)\]

OpenStudy (anonymous):

thx

OpenStudy (chrisplusian):

It can be x that is attached to the 2pi/3 or a t or whatever parameter you use

OpenStudy (baru):

2pi/3 is the period, you have to change it to "frequency"

OpenStudy (chrisplusian):

Do you understand why? Or did you just need to check your answer?

OpenStudy (anonymous):

no

OpenStudy (chrisplusian):

It depends on what context.... the trig books show this as "angular frequency" in radians per second. If you want frequency in Hz for an engineering class then you would need to convert this into Hz

OpenStudy (anonymous):

i dont have a book

OpenStudy (baru):

they given value 2pi/3 is "time period". angular frequence = 2pi/(time period) so angular frequency is "3"

OpenStudy (baru):

y=6sin(3(x+1))